### v2

# pasta: Pattern Analysis for Spatial Omics Data

Abstract
Spatial omics assays allow for the molecular characterisation of cells in their spatial context. Notably, the two main technological streams, imaging-based and high-throughput sequencing-based, can give rise to very different data modalities. The characteristics of the two data types are well known in adjacent fields such as spatial statistics as point patterns and lattice data, and there is a wide range of tools available. This paper discusses the application of spatial statistics to spatially-resolved omics data and in particular, discusses various advantages, challenges, and nuances. This work is accompanied by a vignette, pasta (https://robinsonlabuzh.github.io/pasta), that showcases the usefulness of spatial statistics in biology using several R and Python packages.
Introduction
Molecular profiling of cells in organs or tissues can be accomplished either in bulk or at the single-cell level. However, tissue must be dissociated, which may select against certain cell types and results in the loss of the spatial organization of the cells. This limitation is overcome with spatial profiling techniques, which can range from spatial proteomics (e.g., IMC, MIBI-TOF, 4i, CODEX, mIF [giesenHighlyMultiplexedImaging2014, kerenMIBITOFMultiplexedImaging2019, gutMultiplexedProteinMaps2018, blackCODEXMultiplexedTissue2021, tsujikawaQuantitativeMultiplexImmunohistochemistry2017]), to spatial transcriptomics assays based on fluorescence in situ hybridisation (e.g., FISH, MERFISH, CosMx, Xenium [langer-saferImmunologicalMethodMapping1982, chenSpatiallyResolvedHighly2015, janesickHighResolutionMapping2023]) and based on sequencing (e.g., 10x Visium and Visium HD, Slide-seq V1&V2 [stahlVisualizationAnalysisGene2016, oliveiraCharacterizationImmuneCell2024, rodriquesSlideseqScalableTechnology2019, stickelsHighlySensitiveSpatial2021]) to even spatial epigenomics [luSpatiallyResolvedEpigenomic2022]. Moreover, the combination of individual methods enabled spatial multiomic research [vandereykenMethodsApplicationsSinglecell2023]. The technologies and their application in biological research have been the topic of various detailed reviews [aspSpatiallyResolvedTranscriptomes2020, moffittEmergingLandscapeSpatial2022, raoExploringTissueArchitecture2021, mosesMuseumSpatialTranscriptomics2022, pallaSpatialComponentsMolecular2022, bressanDawnSpatialOmics2023].
In this work, we focus on the concepts and application of exploratory statistical approaches for spatial data with a focus on downstream steps; our overview also centres around key differences between data modalities that result from current spatial profiling technologies. In this regard, we complement the ongoing discussion [zormpasMappingTranscriptomeRealizing2023, mosesVoyagerExploratorySinglecell2023] about the challenges in applying spatial statistics to spatial (transcript)omics data.
Most spatial omics assays can be classified as either high-throughput sequencing (HTS) based or imaging-based. In HTS-based approaches, positional information is recorded using a predetermined array of spots or beads. Imaging-based approaches, however, either target the molecules of interest with fluorescent probes, ablate regions stained with a cocktail of antibodies via metal tag readouts, or amplify and sequence target sequences in situ. Several technologies are emerging, but the main trade-offs stem from the resolution, number of features, and sensitivity of the readout [raoExploringTissueArchitecture2021, parkSpatialOmicsTechnologies2022]. For example, HTS-based approaches capture the entire transcriptome (i.e., untargeted), but come at a resolution determined by the spot size. On the other hand, imaging-based approaches typically have lower depth of information (i.e., targeted) but show a higher resolution in comparison to HTS-based approaches. Note also that the landscape of spatially-resolved molecular technologies is rapidly changing, mostly in terms of higher resolution for HTS-based methods and comprehensive marker panels for imaging-based approaches (Figure 1, Technology) [moffittEmergingLandscapeSpatial2022, raoExploringTissueArchitecture2021, rademacherComparisonSpatialTranscriptomics2024].
In terms of data analysis, the technological streams are quite distinct. HTS-based approaches collect data along regularly-spaced spots or beads resulting in a so-called regular lattice. In contrast, imaging-based approaches measure features at exact locations that can be assumed to have originated from a stochastic process known as a point process [raoExploringTissueArchitecture2021]. Thus, we will distinguish between lattice-based and point pattern-based spatial omic data. While the focus is often on the technology (c.f., imaging-based versus HTS-based), we argue that the distinction should lie at the data representation level [raoExploringTissueArchitecture2021]. This is important, since there are technologies that can be represented as both point patterns and lattice data, depending on the resolution and data processing. For example, imaging-based technologies that detect transcripts at subcellular resolution result in data where transcript locations can be represented as a point pattern, while for the same dataset, segmented cells can be represented either as an irregular lattice, or as a point pattern via their cell centroids. In general, if events of a point pattern are aggregated in specified regions (e.g., transcripts per cell from segmentation), we end up with lattice data (Figure 1, Data modality) [daleSpatialAnalysisGuide2014, pebesmaSpatialDataScience2023].
Since we can represent some spatial omic datasets as either lattice or point pattern, it is important to further understand the assumptions and analysis strategies of such data modalities [cressieStatisticsSpatialData1993]. On the one hand, spatial point pattern analysis assumes that the point locations were generated by a stochastic process, a so-called event-based view of the data; the goal is to study the properties of this process [baddeleySpatialPointPatterns2015]. On the other hand, one could assume that the locations are fixed and known at the time of sampling and study the associated features at each location via an observation-based view of the data, while recognising that the observations in the lattice are not independent due to the spatial structure [anselinWhatSpecialSpatial1989].
Spatial data modalities have been present in other fields for decades, such as geography, and in particular, the field of spatial statistics offers a large analysis toolbox. For example, spatial omics data collected across a lattice can be analysed by exploratory tools, such as Local Moran’s and the bivariate Lee’s statistic. Imaging data generated by a stochastic process can be input for methods of point pattern analysis, including summaries such as Ripley’s or empty-space functions [zuurAnalysingEcologicalData2007, daleSpatialAnalysisGuide2014, baddeleySpatialPointPatterns2015, pebesmaSpatialDataScience2023] (Figure 1, Analysis).
There are already several tools for omics analysis that employ spatial statistics approaches. Amongst others, Voyager is a framework and a collection of use cases for lattice data, written in both R and Python [mosesVoyagerExploratorySinglecell2023]; it provides data structures and functionalities to compute spatial statistics in molecular biology. Voyager comes with comprehensive vignettes for several spatial profiling technologies. The methodological focus of their vignettes is on lattice data analysis and the spdep geospatial package. Since Voyager offers efficient implementations for lattice data analysis in R, we build on their framework in our vignettes for the lattice data component. Other methods written in R include: SPIAT, which contains various tools for spatial analysis, such as: neighbourhood analysis, local metrics and heterogeneity scores etc. [fengSpatialAnalysisSPIAT2023]; scFeatures, a toolbox comprising lattice data analysis functionality such as Moran’s and point pattern metrics as the -function [caoScFeaturesMultiviewRepresentations2022]; spicyR, which uses a compressed version of the global -function for cross-sample comparison [caneteSpicyRSpatialAnalysis2022]; lisaClust, a spatial domain detection method via LISA (local indicators of spatial association) curves [patrickSpatialAnalysisHighly2023]; mxfda, that uses point pattern summary functions for survival analysis [wrobelMxfdaComprehensiveToolkit2024]; spatialDM uses global and local bivariate Moran’s to score colocalisation of ligand-receptor pairs [liSpatialDMRapidIdentification2023]; and MERINGUE, which uses nearest-neighbour autocorrelation for spatially variable gene selection [millerCharacterizingSpatialGene2021]. Methods in Python include squidpy, a package with various spatial statistics tools for both point pattern and lattice data [pallaSquidpyScalableFramework2022] and PySAL a general Python library for spatial analysis [reyPySALPythonLibrary2010]. For a detailed view on general methods development in the field of spatial transcriptomics, we refer readers to the online Supplementary table of the museum of spatial transcriptomics paper [mosesMuseumSpatialTranscriptomics2022].
Here, we will explore the application of spatial statistics to spatially-resolved omics data guided by the two main streams of data, point pattern-based and lattice-based methods. We will compare the two streams and show their strengths and limitations. Finally, we give an outlook on the challenges and research gaps in the field of spatial data analysis for omics data. Furthermore, we provide a vignette (https://robinsonlabuzh.github.io/pasta) that showcases analysis of data from multiple technologies, where concepts and assumptions are discussed in detail, with inline R and Python code. Overall, our resource Pattern Analysis for SpaTial omics datA (pasta) will highlight the usefulness and transferability of existing exploratory spatial statistics approaches in the context of spatial tissue profiling.
Results
Point Pattern Analysis
Definitions
Most imaging-based technologies give a high-resolution readout of either subcellular compartments (e.g., laser ablation pixels in IMC [giesenHighlyMultiplexedImaging2014]) or even individual transcripts (e.g., FISH [feminoVisualizationSingleRNA1998]). In molecule-level technologies, the expression of features (transcript quantifications, ion counts, fluorescent intensities) is recorded where they occur [aspSpatiallyResolvedTranscriptomes2020]. Often, we are interested in the distribution and composition of cell types. In order to annotate cells, we require (good) cell segmentation boundaries (an active area of research not discussed here; see [wangCellSegmentationImage2019, stringerCellposeGeneralistAlgorithm2021, greenwaldWholecellSegmentationTissue2022, petukhovCellSegmentationImagingbased2022, architSegmentAnythingMicroscopy2023]), allowing transcripts to be allocated to a given cell [fuBIDCellBiologicallyinformedSelfsupervised2024], and then cell type inference. Since the locations of the annotated cells are generated by a stochastic biological process, we can approximate e.g., the cell centroids as points and analyse their location as point patterns.
In the analysis of these point patterns, the goal is to make inferences on the point process that generated the data, not on the patterns themselves [baddeleySpatialPointPatterns2015, p. 127]. Points in a point pattern can also carry a mark, which is a (univariate or multivariate) variable associated with a point [baddeleySpatialPointPatterns2015, pp. 147, 563]. In cell biology, we could imagine both categorical (e.g., cell type or spatial domain labels) or continuous marks (e.g., gene expression).
When a point pattern with many cell types is considered, there are two approaches to formulate the setup. In the first variant, we consider cells in a tissue to depend on each other so their distribution is due to one overarching biological process; this is referred to as a multitype view. In contrast, when we consider the patterns to be created by point processes, we assume that these processes can be individually analysed; this represents the multivariable viewpoint (change in terminology in comparison to [baddeleySpatialPointPatterns2015]). The processes do not need to be independent, but the independence of components (and functions) is often used as a null model for multivariable analysis (see box Assumptions) [baddeleySpatialPointPatterns2015, pp. 565ff.].
Assumptions
Many imaging platforms measure patterns in a field of view (FOV) and collect several such FOVs per sample. In some technologies, there are no gaps between the individual FOVs and they can be assembled into one big image. If there are gaps, consecutive FOVs can either be stitched together (which can be computationally complex) or analysed individually [aspSpatiallyResolvedTranscriptomes2020, mohammadiFastRobustFeaturebased2024]. When analysing individual FOVs, we do not observe the entire pattern but rather a subset of the bigger pattern. This is called window sampling because the window is just a sample of the bigger point pattern. A related concept is the small world model, which describes that points can only be observed in a finite world and not beyond these boundaries [baddeleySpatialPointPatterns2015, pp. 143ff.]. Often, the distinction between the window sampling and small world model concepts is not clear. An example is the arrangement of the epithelium, which can can be imaged with several FOVs (window sampling) but there are no cells expected outside of the epithelial boundary (small world) [garrido-trigoMacrophageNeutrophilHeterogeneity2023]. Therefore, in cell biology, we often encounter a mixture of these two concepts.
Apart from the window of measurement, we need to make assumptions on the statistical properties of the point pattern, including, most importantly, whether points can be considered homogeneous or not. Homogeneity (see Box Assumptions) assumes that the number of points in a given region is proportional to its area , for arbitrary ; i.e., homogeneity refers to a uniform intensity of points across the window of measurement. If that is not the case, the point process is said to be inhomogeneous [baddeleySpatialPointPatterns2015, p.132-133]. This difference has important implications for the analysis of a process. For example, when adjusting for varying local intensities, the process could be found to be inhomogeneous rather than genuinely clustered. This is called the confounding between intensity and interaction (see box Assumptions) [baddeleySpatialPointPatterns2015, pp.151-152].
Options for analysis
Point pattern analysis offers a broad spectrum of options. Most often, we compare patterns in 2D (e.g., a single histological slice), but the methods are applicable with some adjustments to 3D patterns (e.g., confocal microscopy volumes) [baddeleySpatialPointPatterns2015, p. 651], with some increased computational cost [baddeleyAnalysisThreeDimensionalPoint1993].
As mentioned above, points can carry marks; the nature of these marks determines what options are available to the analyst. For example, many point pattern methods are designed for categorical data, such as cell types or spatial domain labels [baddeleySpatialPointPatterns2015, p.147] and can be adapted to the analysis of continuous marks [baddeleySpatialPointPatterns2015, pp.644 ff.]. Model-based approaches for spatial data might also be of interest. This is an entire subfield on its own with many options available [baddeleySpatialPointPatterns2015, pp.297 ff.][pebesmaSpatialDataScience2023].
When comparing categorical marks, the analyst must decide whether to compare a single type (within-type comparison) or between a pair of types (cross-type comparison). For example, a within-type comparison could use the pattern of epithelial cells in a microscopy image; the correlation, spacing or distances between epithelial cells could be summarised. In contrast, a cross-type analysis could consider the correlation between epithelial cells and B cells [baddeleySpatialPointPatterns2015, pp. 561ff.].
Examples
A well-known point pattern summary is Ripley’s -function, defined as the expected number of points falling in a specified radius around an arbitrary point, averaged over all points in the observation window (e.g., FOV) (potentially adjusted for edge effects; see box Assumptions). Ripley’s -function thus quantifies the correlation structure of a point pattern, with within-type and cross-type versions. A point pattern where points have no preference for location shows complete spatial randomness (CSR). This means that the number of points that fall in an arbitrary region follow a Poisson distribution with a rate parameter that is proportional to the size of the region. A clustered process would have values of the -function larger than the Poisson process at that specific scale; values of smaller than the Poisson process indicate a self-inhibiting process (i.e., where points repel each other) [ripleySecondOrderAnalysisStationary1976, ripleyModellingSpatialPatterns1977][baddeleySpatialPointPatterns2015, pp.203 ff.].
Ripley’s -function can be variance stabilised by taking the square root, known as Besag’s -function [besag1977contribution]. It is a global estimate that measures correlation across an entire observation window and can be used instead of the standard -function [baddeleySpatialPointPatterns2015, p.402], [baddeleySpatstatPackageAnalyzing2005].
Figure 2A-C show point patterns of mature oligodendrocytes over three different serial sections of the same mouse hypothalamic preoptic region (the numbers , , indicate the -axis position of the slices: one slice every 50 m along the anterior-posterior axis [moffittMolecularSpatialFunctional2018]). Figure 2D shows the homogeneous global -function, which quantifies the correlation within each slice and the respective -function for a Poisson process (dashed line). The homogeneous -function indicates strong clustering, even for the slice, however, the inhomogeneous -function, which makes a local adjustment for intensity, indicates a level of homogeneity very similar to a (completely spatially random) Poisson process (Supp. Figure S1A-C vs. D-F). This is an example of intensity vs. interaction confounding (see box Assumptions). If slice should therefore be interpreted as clustered, depends on the assumption of homogeneity for that slice. In addition, there are differences between the inhomogeneity corrections in Figure S3; the corrections are discussed in more detail in the vignette.
Local versions of the various summary functions exist. The LISA (local indicators of spatial association) framework provides local versions for many spatial statistics (including point pattern and lattice data analysis) [anselinLocalIndicatorsSpatial1995]. For example, Figure 2E shows LISA curves for the slice [getisSecondOrderNeighborhoodAnalysis1987, anselinLocalIndicatorsSpatial1995]. That is, for each point, we calculate a local -function. These curves can be analysed with functional principal component analysis (fPCA), which allows to extract the main functional modes of variation [ramsayPrincipalComponentsAnalysis2005]. The idea to perform fPCA on spatial statistics summaries has been used in ecological and spatial omics data [illianPrincipalComponentAnalysis2006, vuFunSpaceFunctionalSpatial2023, vuSPFSpatialFunctional2022]. Figure 2F shows the scores of the first two functional PCs. There are two main clusters, representing either the physically clustered oligodendrocytes or those in low-density regions.
Challenges and limitations
Analysing biological samples formally as point patterns comes with several challenges and limitations. The first and most important is viewing the biological sample as the realisation of a point process. It is generally not easy to define whether the entire pattern of several cell types should be represented as a single point process (multitype analysis). Analysing the cell types individually and considering them as separate point processes might underestimate dependencies among the cells (multivariable analysis) [baddeleySpatialPointPatterns2015, pp. 565 ff.]. If, as in our example in Figure 2, the sections are far apart from each other, it can be justified to analyse the processes in isolation as a multivariable analysis. However, if we consider cells in two adjacent slices, these have probably been formed by one biological process and should not be analysed in isolation.
Obtaining a representative sample of a point pattern is also a challenge. In spatial omics, we are often provided with FOVs that an experimentalist has selected from a much larger region or that stem from the technology itself. Such a sample might have different characteristics compared to the point pattern of the larger region, as FOVs are sometimes selected based on morphological properties (e.g., H&E staining). In cases where we have a small world scenario, we would want to limit our observation window to the full region where the points can occur.
Lattice Data Analysis
Definitions
In contrast to point pattern-based methods, analysts could take the observation-based view and regard the location of cells or spots as a fixed (regular or irregular) lattice and consider the corresponding marker expression at each cell or spot. Several HTS-based spatial transcriptomics technologies produce data on a regular lattice (i.e., evenly spaced spots or beads of uniform size and shape), whereas imaging-based technologies yield irregular lattice structures (i.e., segmentation resulting in variable cell sizes and shapes, and non-uniform spacing) [zuurAnalysingEcologicalData2007, pp. 321ff.] [anselin2024introduction].
Assumptions
Once we have recorded the outline and arrangement of the spatial units in the lattice, we can specify the strength of spatial relationships between each unit. Each pair of units is assigned a weight: the stronger the connection between units, the higher the weight. The collection of all weights between locations and , , forms the weight matrix (see box Options for analysis). The construction of the weight matrix is critical for all downstream analyses as it encodes the spatial relationship between the units in the lattice [zuurAnalysingEcologicalData2007, pp. 321ff.] and there exist various strategies: contiguity-based (i.e., in direct contact), graph-based, distance-based, or higher-order neighbours [getisSpatialWeightsMatrices2009, zuurAnalysingEcologicalData2007, pebesmaSpatialDataScience2023].
Options for analysis
A common exploratory analysis for lattice data is based on the concept of spatial autocorrelation. Spatial autocorrelation measures the degree of association of features, e.g., the expression of a gene, that are assumed to be dependent in space. For each combination of spatial locations, a measure of association is calculated and scaled by the weight of the connection [anselin2024introduction]. This approach considers both the proximity (via the weight matrix) and the characteristics of the locations (via the metric) [zuurAnalysingEcologicalData2007, pp. 327ff.] [pebesmaSpatialDataScience2023, pp. 209ff.] [getisAnalysisSpatialAssociation1992]. Over the years, various association measures have emerged, each specific to an aspect of the data. While many allow for univariate comparisons, in which one variable (e.g., expression of a gene) is compared over multiple locations, others allow for the comparison of two (bivariate) or more (multivariate) features at once [pebesmaSpatialDataScience2023, mosesVoyagerExploratorySinglecell2023]. For most metrics, the features are assumed to be continuous measurements. However, metrics allow for the measurement of spatial autocorrelation of categorical variables (e.g., join count statistics). In addition to the number of features and the type of measurement that can be compared in the association score, there is the to consider: global spatial autocorrelation metrics estimate the average level of spatial autocorrelation across all locations whereas local measures give a statistic at each location. The global statistic can be seen as the weighted sum of its respective local statistics [anselinLocalIndicatorsSpatial1995].
Example
Moran’s is a prominent example of a spatial autocorrelation metric. It can be interpreted as the mean Pearson correlation between values at locations and the mean value of the corresponding neighbours [moranNotesContinuousStochastic1950]. A positive value of Moran’s indicates that locations with similar values are clustered while a negative value indicates the clustering of dissimilar values, giving a measure of spatial heterogeneity. The global value of Moran’s is bounded by and with an expected value under spatial randomness near for large , (). Figure 3A shows evidence for spatial autocorrelation of Nrgn expression in mouse brain tissue; the (global) Moran’s value is (p-value essentially zero). Global measures can be used to identify spatially variable genes when compared to the global autocorrelation values of different genes [weberNnSVGScalableIdentification2023, chenEvaluatingSpatiallyVariable2024]. Furthermore, there is a local version of Moran’s (interpretation is analogous to the global version). Figure 3B shows the estimate of local Moran’s at each measured spot (n.b., the local version is not bounded by and ). Figure 3C shows the corresponding adjusted p-values (testing for non-zero spatial autocorrelation). When interpreting local autocorrelation measures, it is important to consider both the effect size estimates and the significance level. Since the significance level is calculated for each spot separately, it is recommended to adjust for multiple testing (e.g., Benjamini and Hochberg [benjamini_hochberg1995fdr]). Overall, local Moran’s statistics reveal locations that have similar values to their neighbours. However, as can be seen in Figure 3A-C regions with the highest and most significant local Moran’s value lie in the part of the tissue where expression of gene Nrgn is very low amongst neighbours. Notably, the local Moran’s measure is both dependent on the value of the log-transformed counts and the similarity among neighbours (c.f., Supplementary Figure S2A-B). For the analysis of categorical data, there are metrics such as the join count statistics [daleSpatialAnalysisGuide2014, pp. 141ff.]. In essence, the join count statistic calculates the frequency of categories among neighbours and compares this value with a theoretical distribution or permutations of the labels to get a significance score. This can be used to investigate the interactions between cell types in a lattice since the distance can be specified when constructing the neighbourhood graph.
Challenges and limitations
The choice of the weight matrix poses a challenge in the analysis of lattice data, as it might influence downstream analyses and conclusions. A common choice is to give a non-zero weight only to direct neighbours of a spatial unit (i.e., contiguity-based). Neighbours based on graphs or on distance [pebesmaSpatialDataScience2023, pp. 191 ff.] are also used; the open question is which is best suited for spatial omics data. If one uses contiguity-based neighbours in imaging-based spatial transcriptomics, results ultimately depend heavily on the accuracy of the cell segmentation, which is known to be challenging [stringerCellposeGeneralistAlgorithm2021, greenwaldWholecellSegmentationTissue2022, petukhovCellSegmentationImagingbased2022, architSegmentAnythingMicroscopy2023, wangCellSegmentationImage2019]. If neighbours beyond the adjacent ones are used, the scale (e.g., at which cells interact still) needs to be specified. Figure 3C-D show the difference between the local Moran’s calculation when based on contiguity-based neighbours in D, the 10 nearest neighbours in E or neighbours within a 1000 pixel distance () in F. While the overall differences are small, some cells show different local Moran’s values. Smoothing of the local Moran’s values occurs when more neighbours are considered (Supplementary Figure S2C). Supplementary Figure S2A and B further show that for some cells, no contiguous neighbours were found, which results in a zero estimate of the local Moran’s . Since cells do not function in isolation but form complex anatomical structures including extracellular components, an open question is whether analyses improve when weight matrix construction takes anatomical structures or regions into account. Overall, it remains to be investigated how much the construction of the weight matrix influences downstream analyses in spatial omics data.
Short guide to spatial analysis of biological data
The large variety of options for analysis in spatial statistics can be challenging to navigate. A few fundamental steps should help selecting appropriate exploratory analysis. First, analysts should consider their technology and what options they have (i.e., imaging-based or HTS-based approaches). As seen in Figure 1, the data representation leads to either point pattern or lattice data analysis. If both streams are possible, the decision depends on the research question. In point pattern analysis, the process leading to the locations of individual points is the primary interest. It allows us to make statements such as whether a point process exhibits clustering or if two cell types tend to interact more than randomly expected. In lattice data analysis, however, we are not directly interested in the arrangement of points, but rather locations are important to define the spatial relationship between locations (i.e., via the weight matrix). We study the interaction of features, such as gene expression, given their spatial location. For example, if ligand-receptor dynamics are of interest, the weight matrix can be specified in a way that captures spatial associations of gene expression between direct neighbours or at a specified scale.
Second, the observation scale has to be chosen. Sometimes, this is a technological scale such as multiple FOVs in imaging. Figure 4A shows a single FOV with islets from the human pancreas acquired using IMC [damondMapHumanType2019]. An analysis considering the entire FOV will tell us about the processes that gave rise to the spatial distribution of the endocrine pancreas (islets). As we see in Figure 4B, a Ripley’s -function under a homogeneity assumption indicates that in this observation window, endocrine cells are clustered within pancreatic islets. However, there might also be a relevant biological scale where we subset our observation window to the islets themselves (Figure 4D). In this case, Ripley’s -function indicates a homogeneous distribution of the islets cells within the observation window (Figure 4E). Overall, the correct scale needs to be chosen in accordance with the research question, since the interpretation depends on it.
In point pattern analysis, the analyst needs to decide whether (and how) to control for inhomogeneous intensities, leading back to potential confounding of intensity and interaction [baddeleySpatialPointPatterns2015, pp. 151 ff.]. Analysts can look at smoothed density estimates to inspect potential inhomogeneity prior to analysing the pattern more formally. Furthermore, it depends again on the research question and biological domain knowledge whether the assumption of homogeneity is appropriate. While tissues such as the adipose tissue form regular structures that could be assumed to be homogeneous, other tissues such as stratified epithelial structures appear to be more inhomogeneous. As shown in Figure 4, the homogeneous and inhomogeneous variants of the spatial summaries can lead to different interpretations.
A fundamental choice in lattice data analysis is the construction of the weight matrix. As seen in Figure 3A-B, there are differences in the interpretation of results between different weight matrix choices. The appropriate weight matrix design depends on the research question. For instance, if the feature of interest is a ligand that requires direct cell contact, using contiguous neighbours is a reasonable choice. If the feature of interest is a molecule that diffuses to nearby cells (e.g., paracrine signalling), a nearest neighbour or distance-based approach should be considered for constructing the neighbourhood matrix.
Overall, we advise analysts to consider both scale and homogeneity carefully, keeping the biological context in mind and to document the decisions made and their rationale.
Discussion
In this work, we presented multiple options for the exploratory data analysis of spatial omics data. We introduced the foundations of both point pattern and lattice data analysis and showed the usability of both frameworks in accompanying vignettes.
Fundamentally, point pattern and lattice data analysis are two distinct paths to analysing spatial omics data. Analysts should be aware of the assumptions and interpretations that can be drawn from each approach. In lattice data analysis, we regard the spatial coordinates as fixed and focus on the dependence between neighbours, i.e., weight matrices that reflect these neighbour interactions. In contrast, point pattern analysis concentrates on the underlying mechanisms that generate stochastic spatial patterns (i.e., clustered, regular or spaced arrangements). Therefore, the choice of the analysis framework depends not only on the data modalities, but also on the research question [anselin2024introduction, cressieStatisticsSpatialData1993, baddeleySpatialPointPatterns2015, pebesmaSpatialDataScience2023, daleSpatialAnalysisGuide2014]. Meanwhile, there are dualities between the data modalities from spatially-resolved measurements (c.f., Figure 1). Using cell segmentation, imaging data can be approximated as a points via their centroids. Point locations are stochastic and can therefore be regarded as a realisation of a point process and can be studied using point pattern analysis. On the other hand, we can also view segmentations as an irregular lattice (and aggregate expression per cell). Similarly, HTS-based data with subcellular resolution (e.g., Visium HD [oliveiraCharacterizationImmuneCell2024]) can likewise be interpreted as a regular lattice of (sub)-cellular locations or irregular lattice from segmentations [polanskiBin2cellReconstructsCells2024] or the reconstructed cells can then be approximated by their centroids as points and further analysed using point pattern analysis.
We have focused here on spatial statistics and highlighted two main streams of analysis. Another popular way to represent spatial arrangements is spatial graphs [aliGraphCompassSpatialMetrics2024, windhagerEndendWorkflowMultiplexed2023]. The methods described in the section on lattice data are closely connected to spatial graphs, especially in the construction of a weight matrix. Point pattern analysis concepts are also related to spatial graph via edge rules. For example, Ripley’s can be interpreted as a graph problem with a distance threshold edge rule, as explained in [daleSpatialAnalysisGuide2014, p. 136].
One common challenge in all spatial analyses is the question of the correct scale. In Figure 4, we showed that the scale influences the interpretation of the spatial analysis: what is found to be spatially heterogeneous at a one scale can be locally homogeneous at another [turnerEffectsChangingSpatial1989, daleSpatialAnalysisGuide2014]. Thus, the scale should be defined in accordance with the scientific question.
[…]
There are various other limitations when studying spatially-resolved omics data. On the technical side, imperfect sections can lead to artifacts or non-overlapping FOVs, which can introduce missing data. Furthermore, cells are never perfectly arranged within a histological section, thus cells overlapping in the z-axis are to be expected. In terms of analysis, two-dimensional views neglect processes that occur in three dimensions. For example, in lattice data we might neglect an important contiguous neighbour in an adjacent slice. The same applies to point pattern analysis where we consider only processes in two dimensions even though the underlying biological process likely happens in three dimensions. Both lattice data and point pattern analysis extend conceptually to 3D, however with increased computational complexity. Moreover, in point pattern analysis, we often implicitly assume that patterns are invariant to rotations (see box Assumptions). This assumption is not anatomically accurate for organs and tissues with layered structures, such as the brain. There are other limitations arising from the concept of spatial metrics. Lattice data analysis methods were originally designed for rather low dimensional datasets, while spatial omics often deals with thousands of features (e.g., full transcriptome). Similarly, point pattern methods usually only allow for pairwise comparison of features, which can lead to a high number of cross comparisons e.g., when there are a dozen cell types. Moreover, with newer technologies, the size of the biological sample and resolution of measurements are constantly expanding, which can lead to computational performance issues (e.g., to store cell segmentations in memory to construct weight matrices). Since point pattern based methods only store coordinates, they usually scale better in terms of memory requirements.
In our accompanying vignette, we discuss concepts and assumptions in more detail using biological examples with R and Python. For point pattern analysis, we mainly use R due to the lack of comprehensive Python libraries available for point pattern analysis. Since the R package Voyager is an extensive framework and resource for lattice data, we build on it for the lattice data analysis part of our work. In addition, we provide vignettes for point pattern analysis to complement existing discussions [zormpasMappingTranscriptomeRealizing2023, mosesVoyagerExploratorySinglecell2023].
Outlook
The spatial metrics as described in point pattern and lattice data analysis offer interesting ways to quantify spatial patterns in molecular biology. There are still open research questions including the correct application (e.g., scale and homogeneity) of methods to a given biological context, the high dimensional nature of spatial omics data and the effect of preprocessing steps that come upstream of the application of spatial statistics. Furthermore, comparisons of spatial metrics across multiple samples and conditions and parametric spatial models offer interesting options for future work.
Code and data availability
Code and instructions to retrieve data to reproduce figures and results can be found under https://github.com/robinsonlabuzh/pasta-manuscript, code licensed under the GPLv3 terms. The vignette and source code can be accessed via https://robinsonlabuzh.github.io/pasta licensed under CC-BY-4.0 terms for the text and GPLv3 license for any code.
Acknowledgments
We thank all members of the Robinsonlab for constructive feedback on the manuscript and vignette. In particular, we thank Peiying Cai, Alice Driessen, Reto Gerber, Pierre-Luc Germain, Maruša Kodermann, Vladyslav Korobeynyk, Siyuan Luo, Giulia Moro, Emanuel Sonder, Jiayi Wang and David Wissel for their careful reading and input to the vignette. The Python examples in the vignette were developed as part of the SpaceHack 2024 event. We thank the SpaceHack organisers and participants.
Author contributions
ME: Conceptualisation, Methodology, Software (Vignette), Analysis, Visualisation, Writing - Original Draft; SG: Conceptualisation, Methodology, Software (Vignette), Analysis, Visualisation, Writing - Original Draft; HLC: Conceptualisation, Vignette, Writing - Review & Editing; IM: Methodology, Writing - Review & Editing; MK, Software (Vignette), Writing - Review & Editing; RF: Methodology, Writing - Review & Editing; MDR: Conceptualisation, Methodology, Writing - Original Draft, Supervision, Funding acquisition.
Funding
This work was supported by Swiss National Science Foundation (SNSF) project grant 310030_204869 to MDR. MDR acknowledges support from the University Research Priority Program Evolution in Action at the University of Zurich. HLC acknowledges support by SNSF grant number 222136.
Conflict of interest
We declare no conflict of interest.
Supplementary Figures